One litre CO2 is passed through red hot coke. The volume becomes 1.4 litres at same temperature and pressure. the composition of products is:

Ans: 0.6 L CO2 and 0.8 L CO

Grade:Upto college level

7 Answers

Sunil Kumar FP
10 years ago
co2 + c----2co
let x lt of co2 cokvine with coke it will form 2xlt of co.
voluke of co2 left =1-x
total volume =1.4 lt
1-x+ 2x =1.4
x=.4
vopume of co=2×.4=.8 lt
volume of co2=1-.4 lt=.6 lt
thanks and regards
sunil kumar
askIItian faculty
Manav Aggarwal
6 years ago
Co2(g)+co(s)------2co(s)Initial vol. (1L). 0. 0Final vol. 1-x. 2xFinal vol. 1-x+2x=1.4 X=0.4Vol of co2=2×0.4=0.8Vol of co=(1-0.4)=0.6
Depp
6 years ago
Shit happens when this is asked in exams.
Shubham BHARDWAJ
5 years ago
Co2+C= CO this is the unbalanced reaction when carbon dioxide is passed through red hot coke .for solving questions like this one should balance the above mentioned equation and proceed with solving the problems
kauship saha
5 years ago
The reaction involved -co2 + c= 2co1L 2L Given total volume of composition 1.4L Actual volume of composition is 2L therefore, co2 left= actual volume -given volume =2-1.4 =0.6 literNow, co remains=given volume of composition -volume of co2 =1.4-0.8 = 0.6 liter
saqib souran
5 years ago
co2 + c----2co
let x lt of co2 cokvine with coke it will form 2xlt of co.
voluke of co2 left =1-x
total volume =1.4 lt
1-x+ 2x =1.4
x=.4
vopume of co=2×.4=.8 lt
volume of co2=1-.4 lt=.6 lt
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your question below.
 
CO2 + C → 2CO
let x lt of CO2 cokvine with coke it will form 2x lt of CO.
volume of CO2 left =1 – x
total volume =1.4 lt
1 – x + 2x = 1.4
x =.4
vopume of CO = 2×.4=.8 lt
volume of CO2 = 1 – .4 lt = 0.6 lt
 
Hope it helps.
Thanks and regards,
Kushagra